A convenience store recently started to carry a new brand of soft drink. Management is interested in estimating future sales volume to determine whether it should continue to carry the new brand or replace it with another brand. The following table provides the number of cans sold per week. Use both the trend projection with regression and the exponential smoothing​ (let alpha α=0.4with an initial forecast for week 1 of 567​) methods to forecast demand for week 13. Compare these methods by using the mean absolute deviation and mean absolute percent error performance criteria. Does your analysis suggest that sales are trending and if​ so, by how​ much?

Period   1   2   3   4   5   6   7   8   9   10   11   12
Sales   567   637   650   744   676   631   736   707   754   697   677   720
Observation   1   2   3   4   5   6   7   8   9   10   11   12

Obtain the trend projection with regression forecast.

The forecast for week 13 is ______ (Enter your response rounded to the nearest whole​ number.)
Answer #1

I am going solve this by using excel

1 Period x Sales y 567 637 650 744 676 631 736 707 754 697 677 720 4 4 5 6 6 10 10 12 13 14 15 Slope b 16 Intercepta 623.45 1

We know trend projection with regression forecast equation Y = a+ bx

The forecast for week 13 (x=13) = 623.45 + 9.16*13 = 742.53 or 743

The forecast for week 13 = 743